# SiT: Exploring Flow and Diffusion-based Generative Models with Scalable Interpolant Transformers

### v1

@@ Appendix E Additional Implementation Details
We trained all of our models following identical structure and hyperparameters retained from DiT [48]. We used AdamW [34, 43] as optimizer for all models. We use a constant learning rate of $1\times 10^{-4}$ and a batch size of $256$. We used random horizontal flip with probability of $0.5$ in data augmentation. We did not tune the learning rates, decay/warm up schedules, AdamW parameters, nor use any extra data augmentation or gradient clipping during training. Our largest model, SiT-XL, trains at approximately $6.8$ iters/sec on a TPU v4-64 pod following the above configurations. This speed is slightly faster compared to DiT-XL, which trains at  $6.4$ iters/sec under identical settings. We also gather the training speed of other model sizes and summarize them below.
S B L XL
DiT 20.0 19.8 9.3 6.4
SiT 19.7 20.8 9.3 6.8
Table 2: Training speed (iters/sec) across all model sizes. All training speeds are measured on a TPU v4-64 pod. We note that the training speed is largely influenced by the hardware state.
@@
Linear velocity $\mathcal{L}_{\mathrm{v}}$ 1 0 1 4e-2
@@
Appendix F Additional Visual results
Figure 2: Uncurated $512\times 512$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”volcano”(980)
Figure 2: Uncurated $512\times 512$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”volcano”(980)
Figure 3: Uncurated $512\times 512$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”arctic fox”(279)
Figure 4: Uncurated $512\times 512$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”loggerhead turtle”(33)
Figure 4: Uncurated $512\times 512$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”loggerhead turtle”(33)
Figure 5: Uncurated $512\times 512$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”balloon”(417)
Figure 6: Uncurated $512\times 512$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”red panda”(387)
Figure 6: Uncurated $512\times 512$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”red panda”(387)
Figure 7: Uncurated $512\times 512$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”geyser”(974)
Figure 8: Uncurated $256\times 256$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”macaw”(88)
Figure 8: Uncurated $256\times 256$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”macaw”(88)
Figure 9: Uncurated $256\times 256$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”golden retriever”(207)
Figure 10: Uncurated $256\times 256$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ” ice cream”(928)
Figure 10: Uncurated $256\times 256$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ” ice cream”(928)
Figure 11: Uncurated $256\times 256$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”cliff”(972)
Figure 12: Uncurated $256\times 256$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”husky”(250)
Figure 12: Uncurated $256\times 256$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”husky”(250)
Figure 13: Uncurated $256\times 256$ SiT-XL samples.Classifier-free guidance scale = 4.0 Class label = ”valley”(979)

### v2

@@ Appendix 0.E Additional Implementation Details
We trained all of our models following identical structure and hyperparameters retained from DiT [50]. We used AdamW [35, 45] as optimizer for all models. We use a constant learning rate of $1\times 10^{-4}$ and a batch size of $256$. We used random horizontal flip with probability of $0.5$ in data augmentation. We did not tune the learning rates, decay/warm up schedules, AdamW parameters, nor use any extra data augmentation or gradient clipping during training. Our largest model, SiT-XL, trains at approximately $6.8$ iters/sec on a TPU v4-64 pod following the above configurations. This speed is slightly faster compared to DiT-XL, which trains at  $6.4$ iters/sec under identical settings.
@@
LIN velocity $\mathcal{L}_{\mathrm{v}}$ 1 0 1 4e-2
@@
Appendix 0.F Additional Visual results
Figure 7: Uncurated $512\times 512$ SiT-XL samples.
Classifier-free guidance scale = 4.0
Class label = "volcano"(980)
Figure 8: Uncurated $512\times 512$ SiT-XL samples.
Classifier-free guidance scale = 4.0
Class label = "arctic fox"(279)
Figure 9: Uncurated $512\times 512$ SiT-XL samples.
Classifier-free guidance scale = 4.0
Class label = "loggerhead turtle"(33)
Figure 10: Uncurated $512\times 512$ SiT-XL samples.
Classifier-free guidance scale = 4.0
Class label = "balloon"(417)
Figure 11: Uncurated $512\times 512$ SiT-XL samples.
Classifier-free guidance scale = 4.0
Class label = "red panda"(387)
Figure 12: Uncurated $512\times 512$ SiT-XL samples.
Classifier-free guidance scale = 4.0
Class label = "geyser"(974)
Figure 13: Uncurated $256\times 256$ SiT-XL samples.
Classifier-free guidance scale = 4.0
Class label = "macaw"(88)
Figure 14: Uncurated $256\times 256$ SiT-XL samples.
Classifier-free guidance scale = 4.0
Class label = "golden retriever"(207)
Figure 15: Uncurated $256\times 256$ SiT-XL samples.
Classifier-free guidance scale = 4.0
Class label = " ice cream"(928)
Figure 16: Uncurated $256\times 256$ SiT-XL samples.
Classifier-free guidance scale = 4.0
Class label = "cliff"(972)
Figure 17: Uncurated $256\times 256$ SiT-XL samples.
Classifier-free guidance scale = 4.0
Class label = "husky"(250)
Figure 18: Uncurated $256\times 256$ SiT-XL samples.
Classifier-free guidance scale = 4.0
Class label = "valley"(979)